Ten grams of ground beef are added to 90mL of sterile DI water and blended thoroughly (1:10 dilution).

Then a series of 3 ten-fold dilutions is done using sterile DI water as the diluent. After blending 100uL of the last two dilutions are plated. The most dilute plate had 30 colonies on it and the second most dilute plate had 280 colonies. What is the concentration of bacteria in the ground beef sample?
Answer #1

CFU/ml = number of colonies * diluiton factor / volume plated in ml.

For the dilution that gave 30 colonies

Number of colonies = 30

Dilution factor = 10000 times

Volume plated = 100ul or 0.1ml.

CFU/ml = 30 * 10000 / 0.1

30 * 10 * 10000 = 3 * 106 CFU/ml

For the dilution that gave 280 colonies

Number of colonies = 280

Dilution factor = 1000 times

Volume plated = 100ul or 0.1ml.

CFU/ml = 280 * 1000 / 0.1

280 * 10 * 1000 = 2.9 * 106 CFU/ml

Average of both CFU = 2.8*3 / 2 = 2.9 *106 CFU/ml or 2.9 *106 bacteria/ml.

Given that 10g of beef was added to 90 ml sterile water.

So the total number of bacteria present in 10g beef = 2.9 *106 *10 = 2.9 *107 bacteria.

please write to me if you need more help